# Fast Task Planning with Neuro-Symbolic Relaxation

### v1

@@ IV-C Training
The GNN-based importance predictor $s_{\theta}$ is trained using supervision from optimal demonstrations. Each PDDL task $\tau$ is compiled into a scene graph $\mathcal{G}_{\tau}=(\mathcal{V}_{\tau},\mathcal{E}_{\tau})$, where nodes represent entities $o\in\mathcal{O}_{\tau}$ with feature vectors that include a one-hot type encoding, a multi-hot mask of true unary predicates in the initial state, and a goal-indicator bit for each unary predicate. Edges are created for each ordered pair of distinct objects ($o_{i},o_{j}$), with edge features derived from the binary predicates $\mathcal{P}^{(2)}$, including Boolean indicators for their truth value in the initial state and whether they are required in the goal. Entities are labeled as important ($\ell=1$) if they appear in the ground preconditions or effects of any action in the optimal plan; entities explicitly mentioned in the goal are always labeled as important. The model uses a 3-layer GNN architecture with 16-dimensional hidden layers, ReLU activation, and layer normalization, alternating between edge-to-node and node-to-edge message passing for 3 iterations. Final node representations are mapped to importance scores via a linear layer followed by a sigmoid function. The model is trained using the binary cross-entropy loss defined in (5). While this training process mirrors [10], we emphasize a key difference: in our framework, the GNN’s predictions are used only as a first step. Since GNN-based importance scoring is not always reliable, especially under limited supervision or noisy labels, downstream symbolic planners may get stuck if critical entities are misclassified. This motivates our framework, which augments learned importance scores with rule-based rough-plan recovery and symbolic consistency checks to ensure robustness at inference time.
@@ IV-D Inference
To solve the problem (4), we introduce Flax, a neuro-symbolic relaxation framework designed for fast task planning. It combines neural importance prediction with symbolic rough-plan refinement and rule-based expansion, progressively narrowing down the search space while retaining solvability. The pipeline consisting of three inference steps is illustrated in Fig. 3. To aid understanding, we use the example from Fig. 4 throughout the explanation of each step.
@@
Inference Step 1 (Importance‑guided Pruning)  In the inference stage, we first use the learned GNN importance predictor to assign a score $s(o)\!\in\![0,1]$ to every entity $o$. Starting from a high threshold $q\!\leftarrow\!q_{\max}$, we keep only the entities with $s(o)\geq q$, translate the resulting state into a simplified task, and send it into a symbolic planner. If a solution is returned and verified to be valid on the original problem, the problem is solved; otherwise we gradually lower $q$ until a time budget is exhausted. This pruning strategy offers notable computational advantages by reducing the complexity of the planning problem. But unlike [10], which allocates the entire planning budget to iterative pruning, we dedicate only a small fraction of the total time to this step. Our insight is that aggressive pruning quickly reveals a compact, expressive entity subset that supports further refinement. The importance threshold is initialized to $q_{\max}\!=\!0.81$ with a decay factor $\gamma\!=\!0.9$. After each failure, we set $q\!\leftarrow\!\gamma q$ and rebuild the simplified task $\tau_{\mathcal{O}}$ with the entities $\mathcal{O}\!=\!\{o\!\mid\!s(o)\!\geq\!q\}$, yielding a bootstrap entity set $\mathcal{O}_{1}$.
@@
Inference Step 2 (Rough-plan Enhancement)   Step 1 may exclude critical entities, rendering the simplified task unsolvable. To recover important entities efficiently, we solve a rule‑relaxed version of the original task. Because the relaxed task is much simpler, a rough plan $\pi_{r}$ can be obtained significantly faster than the original problem. The key insight behind this step is leveraging domain-specific relaxation rules to quickly identify entities likely essential for solving the original problem from a symbolic perspective. We collect the set of entities involved in $\pi_{r}$ ($\mathcal{O}_{r}=\{o\mid o\text{ appears in }\pi_{r}\}$) and merge it with $\mathcal{O}_{1}$ to obtain $\mathcal{O}_{2}=\mathcal{O}_{1}\cup\mathcal{O}_{r}$. In the MazeNamo domain, Relaxation Rule is to remove all light boxes, which can be described in natural language as: If an object o is a light box located at position p in the initial state, then eliminate o entirely from the domain description and mark p as empty. This can be expressed in the form of a first-order logic rule:
[…]
@@
For example, in Fig. 4, after two iterations in Step 1, the simplified task remains unsolvable when the time budget of Step 1 expires. In the rule-relaxed task, the solution involves position p20, a critical entity ignored in the importance set, whose absence contributed to the simplified task’s unsolvability. Position p20 is therefore added to the last importance set to form the relaxed importance set.
@@
Inference Step 3 (Complementary Rule Expansion)  Finally, we apply symbolic Complementary Rules $\textsc{Comp}(\cdot)$ that enforce local consistency by ensuring that entities’ initial relationships are maintained in the final importance set. The guiding intuition here is that entities connected by fundamental spatial or relational constraints in the initial state should either both be included or both be excluded. Violating these local consistencies typically results in unsolvable tasks or inefficiencies. The entity set $\mathcal{O}_{3}=\textsc{Comp}(\mathcal{O}_{2})$ defines the final simplified task, which is forwarded to the symbolic planner for the remainder of the time budget. In the MazeNamo domain, we use:
@@
By combining GNN-derived importance, rough plan from rule-based relaxation, and complementary rule expansion, our method recovers missing entities critical for plan validity, while keeping the simplified task size manageable.
@@ V-A Implementation Details
Benchmark  We introduce MazeNamoEnv, a MiniGrid environment that generates random maze navigation tasks on an $n\times n$ grid. The outer boundary consists of walls, and each interior cell is independently assigned as a wall (20%), heavy box (10%), light box (15%), or free space (55%). Tasks are symbolically translated into typed PDDL instances $\tau$ mentioned in Section IV-A. Each instance is solved in satisficing mode and categorized into four difficulty levels (easy, medium, hard, and expert) based on time cost. Unsolvable, trivial ($<\!0.1$s), or excessively difficult instances are discarded. For map sizes $n\!\in\!\{10,12,15\}$, we retain 300, 200, 100, and 100 instances per difficulty level, respectively, with evaluation time limits of 5s, 20s, and 40s. Another 200 easy instances with $n\!=\!10$ are used for training. All the optimal plans are generated by Fast Downward in seq-opt-lmcut mode. And all the satisficing symbolic planners in inference are Fast Downward in LAMA-first mode. To demonstrate sim-to-real generalization, the policy is executed in (i) Isaac Sim with a forklift platform and (ii) the real world using a Unitree Go2 quadruped manipulator.
Figure 5: IsaacSim Experiments. (a) Scene abstraction in a robotics environment. Objects and positions are mapped to symbols. (b)(c)(d) IsaacSim tasks involving different actions.
@@
Adaptation to Robotics Environment   We deploy the exact model trained in MiniGrid simulation on both a forklift in Isaac Sim and a Unitree Go2 quadruped equipped with a 6-DoF D1 manipulator without finetuning. First, a geometry-aware state abstraction maps each object to a symbolic position based on its bounding box, transforming a 3D scene into a partial PDDL instance. Second, to avoid free areas blocking valid path searching, we improves graph connectivity by sampling additional “stepping-stone” positions using circle packing until saturation, as shown in Fig. 5 (a). Lastly, we translates high-level symbolic plans into platform-specific low-level skills, enabling deployment on diverse systems.
Figure 6: Real-world Experiments. A Unitree Go2 quadruped robot, equipped with a 6-DoF D1 arm, a Livox LiDAR, and an Intel RealSense camera (left), performs MazeNamo tasks in real-world environments with varied object layouts. In each task, the robot navigates toward the red flag marker while avoiding or manipulating obstacles. Semi-transparent overlays show the robot’s trajectory over time, illustrating the effectiveness of Flax’s plan execution in complex, real-world scenarios.
@@ V-B Baselines
All baselines share the same GNN architecture and weights, ensuring observed differences are due to the entity-selection strategy rather than the importance estimator.
@@ V-E Real-World Evaluation
To further validate Flax’s practical applicability, as shown in Fig. 6, we deployed the trained model on a Unitree Go2 quadruped robot equipped with a Livox Mid-360 LiDAR, a 6-DoF D1 manipulator, and an Intel RealSense D435 camera mounted near the gripper. These real-world trials encompassed subtasks such as picking up and placing bottles, pushing boxes, and navigating through cluttered spaces. For example, in Fig. 6 (b), where all the chairs are seen as unmovable obstacles, the robot must reach a goal position obstructed by a bottle and multiple boxes. Due to task constraints (two boxes cannot be pushed simultaneously and a box cannot be pushed while a bottle is on top), the robot follows a sequential strategy: ① it navigates to the center box, identifies the bottle as a lightweight, graspable object, lifts it using its arm-mounted gripper, and relocates it to free the box; ② pushes the center box toward the goal to create clearance; ③ moves the two left-side boxes to make space for the box currently occupying the goal; ④ pushes the right-side box to access a position suitable for pushing the center box; and ⑤ finally pushes the center box away and reaches the goal position to complete the task. This experiment showcases a comprehensive integration of complex reasoning and low-level action execution, demonstrating the effectiveness of Flax when planning in physical domains. Despite the inherent complexities of real-world deployment, including sensor noise, minor calibration errors, and non-grid layouts, Flax consistently enabled the robot to achieve its high-level goals. Importantly, this was accomplished without any real-world retraining, using the same policy trained entirely in low-fidelity simulation. Videos of these real-world demonstrations are available at https://sairlab.org/flax/.
@@ VI Conclusion
To summarize, we presented Flax, a three-step neuro-symbolic task-planning framework that combines learned importance prediction with symbolic relaxation and expansion. This approach accelerates planning while improving robustness against incomplete task abstractions. Experiments on the MazeNamo benchmark show a 20.82% increase in success rate and a 17.65% reduction in planning time compared to previous methods. The framework successfully generalizes to real-world tasks without additional fine-tuning, demonstrating its practical applicability. Future work will focus on extending this approach to continuous domains, automatic discovering of symbolic rules, and optimizing the integration of symbolic rules with neural components.
@@ Acknowledgement
This work was partially supported by the DARPA award HR00112490426. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of DARPA.

### v2

@@ IV-C Training
The GNN-based importance predictor $s_{\theta}$ is trained using supervision from optimal demonstrations. Each PDDL task $\tau$ is compiled into a relational graph $\mathcal{G}_{\tau}=(\mathcal{V}_{\tau},\mathcal{E}_{\tau})$, where nodes represent objects $o\in\mathcal{O}_{\tau}$ with feature vectors that include a one-hot type encoding, a multi-hot mask of true unary predicates in the initial state, and a goal-indicator bit for each unary predicate. Edges are created for each ordered pair of distinct objects ($o_{i},o_{j}$), with edge features derived from the binary predicates $\mathcal{P}^{(2)}$, including Boolean indicators for their truth value in the initial state and whether they are required in the goal. Objects are labeled as important ($\ell=1$) if they appear in the grounded preconditions or effects of any action in the optimal plan; objects explicitly mentioned in the goal are always labeled as important. The model uses a 3-layer GNN architecture with 16-dimensional hidden layers, ReLU activation, and layer normalization, alternating between edge-to-node and node-to-edge message passing for 3 iterations. Final node representations are mapped to importance scores via a linear layer followed by a sigmoid function. The model is trained using the binary cross-entropy loss defined in (5). While this training process mirrors [31], we emphasize a key difference: in our framework, the GNN’s predictions are used only as a first step. Since GNN-based importance scoring is not always reliable, especially under limited supervision, downstream symbolic planners may get stuck if critical objects are misclassified. This motivates our framework, which augments learned importance scores with rough-plan recovery and symbolic consistency checks to ensure robustness at planning.
@@ IV-D Planning
To solve the problem (4), we introduce Flax, a neuro-symbolic relaxation framework designed for fast task planning. It combines neural importance prediction with symbolic rough-plan refinement and rule-based expansion, progressively narrowing down the search space while retaining solvability. The pipeline consisting of three planning steps is illustrated in Fig. 3. To aid understanding, we use the example from Fig. 4 throughout the explanation of each step.
@@
Planning Step 1 (Importance‑guided Pruning)  In the planning stage, we first use the learned GNN importance predictor to assign a score $s(o)\!\in\![0,1]$ to every object $o$. Starting from a high threshold $q\!\leftarrow\!q_{\max}$, we keep only the objects with $s(o)\geq q$, translate the resulting state into a simplified task, and send it into a symbolic planner. If a solution is returned and verified to be valid on the original problem, the problem is solved; otherwise we gradually lower $q$ until a time budget is exhausted. This pruning strategy offers notable computational advantages by reducing the complexity of the planning problem. But unlike [31], which allocates the entire planning budget to iterative pruning, we dedicate only a small fraction of the total time to this step. Our insight is that aggressive pruning quickly reveals a compact, expressive object subset that supports further refinement. Threshold is initialized to $q_{\max}\!=\!0.81$ with a decay factor $\gamma\!=\!0.9$. After each failure, we set $q\!\leftarrow\!\gamma q$ and rebuild the simplified task $\tau_{\mathcal{O}}$ with $\mathcal{O}\!=\!\{o\!\mid\!s(o)\!\geq\!q\}$, yielding a bootstrap object set $\mathcal{O}_{1}$.
@@
Planning Step 2 (Rough-plan Enhancement)   Step 1 may exclude critical objects, rendering the simplified task unsolvable. To recover important objects efficiently, we solve a rule‑relaxed version of the original task. The design of the relaxation rules follows a simple guideline: making the problem substantially easier while still preserving a rough structural resemblance to the original task. Because the relaxed task is much simpler, a rough plan $\mu_{r}$ can be obtained significantly faster than the original problem. The key insight behind this step is leveraging domain-specific relaxation rules to quickly identify objects likely essential for solving the original problem from a symbolic perspective. We collect the set of objects involved in $\mu_{r}$ ($\mathcal{O}_{r}=\{o\mid o\text{ appears in }\mu_{r}\}$) and merge it with $\mathcal{O}_{1}$ to obtain $\mathcal{O}_{2}=\mathcal{O}_{1}\cup\mathcal{O}_{r}$. In the MazeNamo domain, Relaxation Rule is to remove all light boxes. This can be expressed as a first-order logic rule:
$\forall\texttt{o}\,\forall\texttt{p}\;\Bigl(\texttt{isObstacle(o)}\land\texttt{isPos(p)}\land\texttt{isMovable(o)}$
$\qquad\quad\land\texttt{isLight(o)}\land\texttt{oAt(o,p)}$
$\;\;\rightarrow\;\;\neg\texttt{isObstacle(o)}\land\neg\texttt{oAt(o,p)}\land\neg\texttt{isLight(o)}$
$\qquad\quad\land\neg\texttt{isMovable(o)}\land\texttt{isEmpty(p)}\Bigr).$
@@
For example, in Fig. 4, after two iterations in Step 1, the simplified task remains unsolvable when the time budget of Step 1 expires. In the rule-relaxed task, the solution involves position p20, a critical object ignored in the importance set, whose absence contributed to the simplified task’s unsolvability. Position p20 is therefore added to the last importance set to form the relaxed importance set.
@@
Planning Step 3 (Complementary Rule Expansion)  Finally, we apply symbolic Complementary Rules $\textsc{Comp}(\cdot)$ that enforce local consistency by ensuring that objects’ initial relationships are maintained in the final importance set. The guiding intuition here is that objects connected by fundamental spatial or relational constraints in the initial state should either both be included or both be excluded. Violating these local consistencies typically results in unsolvable tasks or inefficiencies. The object set $\mathcal{O}_{3}=\textsc{Comp}(\mathcal{O}_{2})$ defines the final simplified task, which is forwarded to the symbolic planner. In the MazeNamo domain, we use:
@@
By combining GNN-derived importance, rough plan from rule-based relaxation, and complementary rule expansion, our method recovers missing objects critical for plan validity, while keeping the simplified task size manageable.
@@ V-A Implementation Details
Benchmark  We introduce MazeNamoEnv, a MiniGrid environment that generates random maze navigation tasks on an $n\times n$ grid. The outer boundary consists of walls, and each interior cell is independently assigned as a wall (20%), heavy box (10%), light box (15%), or free space (55%). Tasks are symbolically translated into typed PDDL instances $\tau$ mentioned in Section IV-A. Each instance is solved in satisficing mode and categorized into four difficulty levels (easy, medium, hard, and expert) based on time cost. Unsolvable, trivial (<0.1<\!0.1s), or excessively difficult instances are discarded. For map sizes $n\!\in\!\{10,12,15\}$, we retain 300, 200, 100, and 100 instances per difficulty level, respectively, with evaluation time limits of 5s, 20s, and 40s. Another 200 easy instances with n=10n\!=\!10 are used for training. All the optimal plans are generated by Fast Downward in seq-opt-lmcut mode. And all the satisficing symbolic planners in planning are Fast Downward in LAMA-first mode. To demonstrate sim-to-real generalization, the policy is executed in (i) Isaac Sim with a forklift platform and (ii) the real world using a Unitree Go2 quadruped manipulator with a D455 camera on the gripper.
Figure 5: IsaacSim Experiments. (a) Scene abstraction in a robotics environment. Obstacles and positions are mapped to symbols. (b)(c)(d) IsaacSim tasks involving different actions.
@@
Adaptation to Robotics Environment   We deploy the exact model trained in MiniGrid simulation on both a forklift in Isaac Sim and a Unitree Go2 quadruped equipped with a 6-DoF D1 manipulator without finetuning. First, a geometry-aware state abstraction maps each obstacle to a symbolic position based on its bounding box, transforming a 3D scene into a partial PDDL instance. Second, to avoid free areas blocking valid path searching, we improves graph connectivity by sampling additional “stepping-stone” positions using circle packing until saturation, as shown in Fig. 5 (a). Lastly, we translates high-level symbolic plans into platform-specific low-level skills, enabling deployment on diverse systems.
Figure 6: Real-world Experiments. A Unitree Go2 quadruped robot, equipped with a 6-DoF D1 arm, a Livox LiDAR, and an Intel RealSense camera (left), performs MazeNamo tasks in real-world environments with varied layouts. In each task, the robot navigates toward the red flag marker while avoiding or manipulating obstacles. Semi-transparent overlays show the robot’s trajectory over time, illustrating the effectiveness of Flax’s plan execution in complex, real-world scenarios.
@@ V-B Baselines
All baselines share the same GNN architecture and weights, ensuring observed differences are due to the object-selection strategy rather than the importance estimator.
@@ V-E Real-World Evaluation
To further validate Flax’s practical applicability, as shown in Fig. 6, we deployed the trained model on a Unitree Go2 quadruped robot equipped with a Livox Mid-360 LiDAR, a 6-DoF D1 manipulator, and an Intel RealSense D455 camera mounted near the gripper. We evaluated six real-world tasks, each executed three times. These real-world trials encompassed subtasks such as picking up and placing bottles, pushing boxes, and navigating through cluttered spaces. For example, in Fig. 6 (b) (Task 4), where all the chairs are seen as unmovable obstacles, the robot must reach a goal position obstructed by a bottle and multiple boxes. Due to task constraints (two boxes cannot be pushed simultaneously and a box cannot be pushed while a bottle is on top), the robot follows a sequential strategy: ① it navigates to the center box, lifts the bottle, and relocates it to free the box; ② pushes the center box toward the goal to create clearance; ③ moves the two left-side boxes to make space for the box currently occupying the goal; ④ pushes the right-side box to access a position suitable for pushing the center box; and ⑤ finally pushes the center box away and reaches the goal position to complete the task. Table II reports the success rates. Our system achieved an average success rate of 67%, despite significant real-world challenges including inaccurate object localization and imperfect physical interactions.
TABLE II: Real-world evaluation on Go2 with a D1 arm.
Task ID Success Rate Failure Causes
1 3/3 — (No failures)
2 2/3 Box rotation causing path blockage
3 2/3 Box rotation causing path blockage
4 1/3 Box rotation causing path blockage
5 2/3 Bottle grasping failure
6 2/3 Robot stuck when turning in a corner
Average 67%
The most common failure mode (Tasks 2–4) arises from the Go2 robot pushing boxes using its head, which provides only a single-point contact. When the push is not perfectly centered, the box may unexpectedly rotate, blocking the intended path. Additional failures include occasional bottle grasping errors (Task 5) and the robot becoming stuck during tight corner turns (Task 6). This experiment showcases a comprehensive integration of complex reasoning and low-level action execution, demonstrating the effectiveness of Flax when planning in physical domains. Despite the inherent complexities of real-world deployment, including sensor noise, minor calibration errors, and non-grid layouts, Flax consistently enabled the robot to achieve its high-level goals. Importantly, this was accomplished without any real-world retraining, using the same policy trained entirely in low-fidelity simulation. Source code and videos of these real-world demonstrations are available at https://sairlab.org/flax/.
@@ VI Conclusion
To summarize, we presented Flax, a three-step neuro-symbolic task-planning framework that combines learned importance prediction with symbolic relaxation and expansion. This approach accelerates planning while improving robustness against incomplete task abstractions. Experiments on the MazeNamo benchmark show a 20.82% increase in success rate and a 17.65% reduction in planning time compared to previous methods. The framework successfully generalizes to real-world tasks without additional fine-tuning, demonstrating its practical applicability. While this work focuses on offline bi-level training and the runtime relaxation–complementation pipeline, the broader bi-level perspective suggests several promising directions, including online refinement of the importance predictor, self-supervised incorporation of plans generated during deployment, and tighter integration between symbolic relaxations and learned relational models. Extending Flax to continuous/hybrid task-and-motion domains, as well as enabling automatic discovery of relaxation rules, also represents valuable future work.
@@ Acknowledgement
Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of DARPA or Sony.